OpenStudy (anonymous):

Marvin currently has a balance of $1,192 in an account he has held for 14 years. He opened the account with an initial deposit of $800. What is the simple interest on the account?

OpenStudy (anonymous):

x=.029

OpenStudy (anonymous):

I didn't really do it the smart way. I just set up the interest equation, which is 1,192=800(1+x)^14 and guess and tested :/

OpenStudy (anonymous):

after solving down to (1+x)^14=1.49, it was kind of easy.

OpenStudy (mathstudent55):

If he deposited $800 and now has $1,192, he earned $392 in interest. Since he earned $392 in interest in 14 years, he earned $392/14 = $28 per year. $28 is i percent of $800? 28 = (i/100)800 i = 3.5%